OpenStudy (anonymous):

Evaluate 7 + (-4x2) for x = 0.

OpenStudy (anonymous):

substitute 0 for x and solve

OpenStudy (anonymous):

\[7+(-4*0^2)=7+(-4*0)=7+0=7\]

OpenStudy (anonymous):

7

OpenStudy (liv1234):

Remember PEMDAS (Parentheses, Exponents, Multiplication, Division, Addition, Subtraction) -4*0=0*2=0. 7+0=7. Your answer is going to be 7.